# Teaching LLMs to See and Guide: Context-Aware Real-Time Assistance in Augmented Reality

### v2

@@ Abstract
The growing adoption of augmented and virtual reality (AR/VR) technologies in industrial training and on-the-job assistance has created new opportunities for intelligent, context-aware support systems. As workers perform complex tasks guided by AR/VR, these devices capture rich streams of multimodal data—including gaze, hand actions, and task progression—that can reveal user intent and task state in real time. Leveraging this information effectively remains a major challenge. In this work, we present a context-aware large language model (LLM) assistant that integrates diverse data modalities, such as hand actions, task steps, and dialogue history, into a unified framework for real-time question answering. To systematically study how context influences performance, we introduce an incremental prompting framework, where each model version receives progressively richer contextual inputs. Using the HoloAssist dataset, which records AR-guided task executions, we evaluate how each modality contributes to the assistant’s effectiveness. Our experiments show that incorporating multimodal context significantly improves the accuracy and relevance of responses. These findings highlight the potential of LLM-driven multimodal integration to enable adaptive, intuitive assistance for AR/VR-based industrial training and assistance.
@@ I Introduction
  • We introduce an incremental prompting framework that systematically evaluates how progressively richer contextual inputs influence model performance. This setup enables a systematic evaluation of how varying levels of context affect the model’s ability to provide accurate guidance and respond to user queries. Additionally, we conduct two ablation studies to assess the individual and combined contributions of each modality, offering practical insights for modality selection under resource constraints.
@@
Section II reviews related literature on AR-based task guidance, multimodal learning, and LLM-assisted interaction systems. Section III presents the proposed context-aware framework, including its multimodal data integration and incremental prompting design. Section IV describes the experimental setup and outlines the evaluation procedures. Section V reports the experimental findings, while Section VI discusses their broader implications and limitations. Section VII concludes the paper and outlines directions for future research.
@@ II RELATED WORK
This section surveys related work on AR task guidance, multimodal context modeling, and the integration of LLMs for adaptive, real-time assistance.
@@ III-A Models
To systematically evaluate the effect of contextual richness on model performance, we designed four incremental model configurations, each incorporating progressively more detailed multimodal inputs. The following summarizes their configurations and intended purposes (see Figure 1).
@@ III-B2 LLM-as-a-judge
To overcome lexical comparison limitations, we adopt the increasingly popular paradigm of using a LLM as a judge. This approach, widely used in recent benchmarks and model evaluations [50, 47, 49, 51, 52], leverages an external LLM (e.g., GPT-4) to assess the quality of generated responses across multiple criteria such as correctness, completeness, relevance, and clarity. As noted by [50], LLM-as-a-judge frameworks have been applied across various domains, including text evaluation, model comparison, and decision-support tasks. Moreover, they offer a cost-effective and consistent alternative to large-scale human evaluations, reducing variability caused by subjective human biases. Despite their advantages, LLM-based evaluators also come with inherent limitations. Recent studies have identified issues such as positional bias, where the model tends to favor the first or longer response, as well as verbosity bias, self-consistency drift, and sensitivity to prompt phrasing [52, 47, 48]. Moreover, because the judging model often shares similar pretraining data and reasoning patterns with the evaluated models, it may reinforce shared misconceptions or stylistic artifacts, leading to overestimation of response quality. These factors underscore the importance of human validation and careful prompt design when interpreting LLM-as-a-judge results.
@@ IV-A HoloAssist Dataset
In more detail, we consider sixteen types of tasks, each of which contains more than two user questions, making them strong candidates for evaluating the LLM assistant performance. Overall, we work with 169 task samples and a total of 625 questions. Ground truth responses are available for each question, allowing for quantitative comparison and evaluation of context-aware LLMs-generated responses across varying levels of contextual awareness. To effectively leverage the multimodal data provided by HoloAssist, we extract and process a set of contextual components that represent different levels of task information granularity used in our incremental prompting framework.
@@ IV-B1 Task Description
The task description represents the most fundamental layer of contextual information available while a user performs a task. It outlines the general sequence of steps typically followed from start to completion, along with an approximate estimate of the time required to complete the task. Although the HoloAssist dataset does not include a dedicated annotation for task descriptions, it has Narration annotation with a “Long-form Description” attribute. In this field, annotators were instructed to watch the videos and write descriptive paragraphs summarizing the observed activities. To construct a unified task description for each task type, we leveraged the Narration annotations across multiple video samples. Specifically, we employed GPT-4 to aggregate and synthesize the common procedural steps shared among different instances of the same task. This process produced concise descriptions that capture both the overall workflow and the estimated duration of the task. The prompt used to extract task descriptions and approximate durations is provided in Appendix A.
@@ IV-D Prompt Design and API Configuration
All models were implemented using the Azure OpenAI GPT-4 REST API through a Python interface, following a standardized configuration to ensure consistent evaluation across experiments. Each API call contained structured messages composed of a system instruction and a user query. To prioritize accuracy and consistency over creativity, the temperature was set to 0.7. This configuration maintains stable responses with limited randomness, reducing stylistic variability while preserving contextual precision. Since instructor responses in the dataset are typically short and concise, the maximum token length was limited to 100. Each prompt consisted of two components: (1) a system instruction defining the model’s role as a task guidance assistant, and (2) a user input including the user’s question along with all contextual information available up to the current time step ($T$). Higher-level models received progressively richer context, as summarized in Table I. Full prompt templates for all four models are provided in Appendix B. All model outputs were programmatically logged and subsequently evaluated using a combination of lexical comparison, LLM-as-a-judge, and human evaluation to comprehensively assess responses quality.
@@ V-B LLM-as-a-Judge
To complement the quantitative metrics, we employed a LLM as an automatic judge to assess response quality more holistically. The judge model was provided with all contextual information available up to the point each question was asked, including task description, current task step, hand actions data, and conversation history, and was instructed to evaluate the four model-generated responses based on correctness, completeness, contextual relevance, and clarity. This LLM-based evaluation enables the capture of semantic and pragmatic aspects that lexical similarity metrics cannot reliably measure. To mitigate potential positional bias in the judging process (i.e., the tendency of an LLM to favor responses appearing earlier in the list), we randomized the order of the four model outputs across four different permutations: 1234, 3421, 4321, and 2143. While exhaustively evaluating all $4!$ orderings would be computationally expensive, we selected four representative permutations that allow each model to appear in every possible position, which we believe sufficiently mitigates positional bias. The final score for each model represents the average across these permutations. Each model received individual 0–5 ratings on all dimensions, along with an overall comparison summary identifying the most contextually appropriate response. The full judging prompt template is provided in Appendix C.
@@ V-B1 Robustness Across LLM Variants
To evaluate the robustness and model-agnostic nature of our LLM-as-a-judge framework, we replicated the evaluation using three different LLM models: GPT-4o, GPT-4.1, and GPT-5. As illustrated in Figure 5, all models exhibit the same monotonic improvement pattern across the four contextual configurations. The average final scores increase steadily from Model 1 to Model 4 for all judges, with only minor numerical differences in absolute scores.
@@ V-C Human Evaluation
Figure 6: Human evaluation results for all models. The mean final score increases consistently from Model 1 to Model 4, with lower variance at higher context levels, indicating improved response quality and stability.
Figure 7: Winner share distribution from human evaluation. Each bar represents the proportion of instances where a model was rated highest by human evaluators. Model 4 dominates with $81.4\%$ of cases, demonstrating the advantage of richer contextual grounding. Percentages exceed $100\%$ due to ties across models.
@@ V-D LLM–Human Agreement
Figure 8: Comparison of LLM-as-a-judge and human evaluation trends. Both follow a similar improvement trajectory across models, confirming the high alignment between automated and human judgment.
@@
Figure 9: Agreement matrix between LLM and human winners. Each cell represents the number of shared wins for a given model pair. An overall agreement rate of $83.7\%$ indicates strong consistency between LLM and human evaluators. Multiple winners per case were included to account for ties.
@@ V-E1 All-Except-One (AEO)
Figure 10: Winner-share comparison across models in the All-Except-One (AEO) setup. The models missing task description or current task step information (Models I and III) show similar declines. The absence of hand action data (Model IV) results in the steepest drop, underscoring its critical role in maintaining robust model performance.
@@ V-E2 Only-One
Figure 12: Winner-share comparison for the Only-One (OO) analysis. Each bar indicates the proportion of cases where a model’s response achieved the highest quality score according to LLM-as-a-Judge evaluation. The hand action context (Model d) attains the highest win rate, reinforcing its importance as the most reliable standalone signal for real-time task understanding.
Figure 13: Boxplot of evaluation scores in the Only-One (OO) configuration, where only one contextual component is provided at a time. The current task step and hand action contexts yield the highest median scores (around 4), while task description also supports strong performance. Model c achieves the slightly higher median, whereas Model d exhibits greater consistency with fewer extreme outliers, indicating more stable behavior across samples.
@@ VI Discussion
In this study, we introduced a comprehensive framework for context-aware, real-time LLM assistants designed to support task guidance in AR environments. The observed improvements in response quality with the addition of contextual information indicate that fine-grained physical and semantic context substantially enhances the performance of LLM-based assistants. While our analysis was conducted using annotated tasks from the HoloAssist dataset—which may not fully capture the variability of real-world scenarios—our findings provide a strong foundation for future deployment and exploration. Yet, several fundamental research directions remain:
@@ VII Conclusions
Appendix A Prompt for Task Description Extraction
2.[Narration 2] ([Duration 2]s)
Appendix B Models’ Prompt Templates

### v3

@@ Abstract
The growing adoption of augmented and virtual reality (AR/VR) technologies in industrial training and on-the-job assistance has created new opportunities for intelligent, context-aware support systems. As workers perform complex tasks guided by AR/VR, these devices capture rich streams of multimodal data—including gaze, hand actions, and task progression—that can reveal user intent and task state in real time. Leveraging this information effectively remains a major challenge. In this work, we present a context-aware LLM assistant that integrates diverse data modalities, such as task description, hand actions, and dialogue history, into a unified framework for real-time question answering. To systematically study how context influences performance, we introduce an incremental prompting framework, where each model version receives progressively richer contextual inputs. Using the HoloAssist dataset, which records AR-guided task executions, we evaluate how each modality contributes to the assistant’s effectiveness. Our experiments show that incorporating multimodal context significantly improves the accuracy and relevance of responses. These findings highlight the potential of LLM-driven multimodal integration to enable adaptive, real-time, and intuitive assistance for AR/VR-based industrial training and assistance.
@@ I Introduction
  • We introduce an incremental prompting framework that systematically evaluates how progressively richer contextual inputs influence model performance. This setup enables a systematic evaluation of how varying levels of contextual information affect the model’s ability to provide accurate guidance and respond to user queries. Additionally, we conduct two ablation studies to assess the individual and combined contributions of each modality, offering practical insights for modality selection under resource constraints.
@@
Section II reviews related literature on AR-based task guidance, multimodal learning, and LLM-assisted interaction systems in AR/VR environments. Section III presents the proposed context-aware framework, including its multimodal data integration and incremental prompting design. Section IV describes the experimental setup and outlines the evaluation procedures. Section V reports the experimental findings, while Section VI discusses their broader implications and limitations. Section VII concludes the paper and outlines directions for future research.
@@ II RELATED WORK
This section surveys related work on multimodal context modeling, LLMs integration for adaptive, real-time assistance, and AR/VR task guidance.
@@ III-A Models
To systematically evaluate the effect of contextual richness on model performance, we designed four incremental model configurations, each incorporating progressively more detailed multimodal inputs. The following summarizes their configurations and intended purposes (see Figure 1). All models were implemented using the Azure OpenAI API (GPT-4o, version: 2025-01-01-preview). Each query was executed with a maximum token limit of 100 to mimic concise, instructor-style responses. For detailed prompt templates of each model configuration, please refer to Appendix A.
@@ III-B2 LLM-as-a-judge
To overcome lexical comparison limitations, we adopt the increasingly popular paradigm of using a LLM as a judge. This approach, widely used in recent benchmarks and model evaluations [50, 47, 49, 51, 52], leverages an external LLM (e.g., GPT-4) to assess the quality of generated responses across multiple criteria such as correctness, completeness, relevance, and clarity. As noted by [50], LLM-as-a-judge frameworks have been applied across various domains, including text evaluation, model comparison, and decision-support tasks. Moreover, they offer a cost-effective and consistent alternative to large-scale human evaluations, reducing variability caused by subjective human biases. Despite their advantages, LLM-based evaluators also come with inherent limitations. Recent studies have identified issues such as positional bias, where the model tends to favor the first or longer response, as well as verbosity bias, self-consistency drift, and sensitivity to prompt phrasing [52, 47, 48]. Moreover, because the judging model often shares similar pretraining data and reasoning patterns with the evaluated models, it may reinforce shared misconceptions or stylistic artifacts, leading to overestimation of response quality. These factors underscore the importance of human validation and careful prompt design when interpreting LLM-as-a-judge results. To mitigate positional bias in our evaluation, we consider the average judgment scores across multiple permutations of model responses order, ensuring that no model is systematically advantaged by its position in the comparison prompt.
@@ IV-A HoloAssist Dataset
In more detail, we consider sixteen types of tasks, each of which contains more than two user questions, making them strong candidates for evaluating the LLM assistant performance. Overall, we work with 141 task samples and a total of 625 questions. Ground truth responses are available for each question, allowing for quantitative comparison and evaluation of context-aware LLMs-generated responses across varying levels of contextual awareness. To effectively leverage the multimodal data provided by HoloAssist, we extract and process a set of contextual components that represent different levels of task information granularity used in our incremental prompting framework.
@@ IV-B1 Task Description
The task description represents the most fundamental layer of contextual information available while a user performs a task. It outlines the general sequence of steps typically followed from start to completion, along with an approximate estimate of the time required to complete the task. Although the HoloAssist dataset does not include a dedicated annotation for task descriptions, it has Narration annotation with a “Long-form Description” attribute. In this field, annotators were instructed to watch the videos and write descriptive paragraphs summarizing the observed activities. To construct a unified task description for each task type, we leveraged the Narration annotations across multiple video samples. Specifically, we employed GPT-4 to aggregate and synthesize the common procedural steps shared among different instances of the same task. This process produced concise descriptions that capture both the overall workflow and the estimated duration of the task. The prompt used to extract task descriptions and approximate durations is provided in Appendix B.
@@ IV-D Prompt Design and API Configuration
All models were implemented using the Azure OpenAI API through a Python interface, following a standardized configuration to ensure consistent evaluation across experiments. Each API call contained structured messages composed of a system instruction and a user query. To prioritize accuracy and consistency over creativity, the temperature was set to 0.5. This configuration maintains stable responses with limited randomness, reducing stylistic variability while preserving contextual precision. Since instructor responses in the dataset are typically short and concise, the maximum token length was limited to 100. Each prompt consisted of three components: (1) a system instruction defining the model’s role as a task guidance assistant, and (2) a user input including the user’s question along with (3) all contextual information available up to the current time step ($T$). Higher-level models received progressively richer context, as summarized in Table I. Full prompt templates for all four models are provided in Appendix A. All model outputs were programmatically logged and subsequently evaluated using a combination of lexical comparison, LLM-as-a-judge, and human evaluation to comprehensively assess responses quality.
@@ V-B LLM-as-a-Judge
To complement the quantitative metrics, we employed a LLM as an automatic judge to assess response quality more holistically. The judge model was provided with all contextual information available up to the point each question was asked, including task description, current task step, hand actions data, and conversation history, and was instructed to evaluate the four model-generated responses based on correctness, completeness, contextual relevance, and clarity. This LLM-based evaluation enables the capture of semantic and pragmatic aspects that lexical similarity metrics cannot reliably measure. To mitigate potential positional bias in the judging process (i.e., the tendency of an LLM to favor responses appearing earlier in the list), we randomized the order of the four model outputs across four different permutations: 1234, 3421, 4321, and 2143, where each number represents the index of a model in the evaluation prompt (e.g., in permutation 3421, Model 3 is presented first, followed by Models 4, 2, and 1). While exhaustively evaluating all $4!$ orderings would be computationally expensive, we selected four representative permutations that allow each model to appear in every possible position, which we believe sufficiently mitigates positional bias. The final score for each model represents the average across these permutations. Each model received individual 0–5 ratings on all dimensions, along with an overall comparison summary identifying the most contextually appropriate response. The full judging prompt template is provided in Appendix C.
@@ V-B1 Robustness Across LLM Variants
To evaluate the robustness and model-agnostic nature of our LLM-as-a-judge framework, we replicated the evaluation using three different LLM models: GPT-4o (version: 2025-01-01-preview), GPT-4.1 (version: 2025-01-01-preview), and GPT-5 (version: 2025-04-01-preview). As illustrated in Figure 5, all models exhibit the same monotonic improvement pattern across the four contextual configurations. The average final scores increase steadily from Model 1 to Model 4 for all judges, with only minor numerical differences in absolute scores.
@@ V-C Human Evaluation
Figure 6: Human evaluation (mean ± standard deviation) results for all models. The mean final score increases consistently from Model 1 to Model 4, with lower variance at higher context levels, indicating improved response quality and stability.
Figure 7: Winner share distribution from human evaluation. Each bar represents the proportion of instances where a model was rated highest by human evaluators. Model 4 dominates with winning $81.4\%$ of cases, demonstrating the advantage of richer contextual grounding. Percentages exceed $100\%$ due to ties across models.
@@ V-D LLM–Human Agreement
Figure 8: Comparison of LLM-as-a-judge and human evaluation trends. Both follow a similar improvement trajectory across models, confirming the high alignment between automated LLM-as-a-judge and human judgment.
@@
Figure 9: Agreement matrix between LLM and human winners. Each cell represents the number of shared wins for a given model pair. An overall agreement rate of $83.7\%$ indicates strong consistency between LLM-as-a-judge and human evaluators. Multiple winners per case were included to account for ties.
@@ V-E1 All-Except-One (AEO)
Figure 10: Winner share comparison across models in the All-Except-One (AEO) setup. The models missing task steps or current task step information (Models II and III) show similar declines. The absence of hand action data (Model IV) results in the steepest drop, underscoring its critical role in maintaining robust model performance.
@@ V-E2 Only-One (OO)
Figure 12: Winner-share comparison for the Only-One (OO) analysis. Each bar indicates the proportion of cases where a model’s response achieved the highest quality score according to LLM-as-a-Judge evaluation. The hand action context (Model d) attains the highest winner share, reinforcing its importance as the most reliable standalone signal for real-time task understanding.
Figure 13: Boxplot of evaluation scores in the Only-One (OO) configuration, where only one contextual component is provided at a time. The current task step (Model c) and hand action (Model d) contexts yield the highest median scores (around 4), while task steps (Model b) also supports strong performance. Model c achieves the slightly higher median, whereas Model d exhibits greater consistency with fewer extreme outliers, indicating more stable behavior across samples.
@@ VI Discussion
In this study, we introduced a comprehensive framework for context-aware, real-time LLM assistants designed to support task guidance in AR environments. Our findings highlight that fine-grained physical and semantic context—captured through multimodal cues such as hand actions, task steps, and dialogue—substantially enhances the reasoning and response quality of LLM-based assistants. This demonstrates the strong ability of LLMs to act as competent, real-time instructors in AR/VR settings, moving beyond static, pre-scripted guidance toward adaptive, interactive task support.
The framework’s primary strength lies in its scalability and generality: it can integrate new sensing modalities and prompting strategies without retraining, enabling flexible deployment across various AR/VR systems. However, its current performance still depends on high-quality contextual annotations and structured task representations, which may limit generalization to unstructured environments. Despite these challenges, the study provides clear evidence that contextual grounding allows LLMs to bridge perception and reasoning, offering practical pathways for intelligent, human-centered task assistance.
Yet, several fundamental research directions remain:
@@
Data Availability
The data used in this study are publicly available. The multimodal dataset employed in this work is based on the HoloAssist dataset released by Microsoft Research [18]. The associated code and sample data for the proposed workflow reproduction are hosted at this GitHub reposito GitHub Repository. Researchers may reuse or extend the dataset and code for further applications and benchmarking, provided that they properly cite this work.
@@ VII Conclusions
Appendix A Models’ Prompt Templates
@@
Appendix B Prompt for Task Description Extraction
2.[Narration 2] ([Duration 2]s)